OpenStudy (anonymous):

A ladder is resting against a wall. The top of the ladder touches the wall at a height of 12 ft. Find the length of the ladder if the length is four ft more than the distance from the wall.

OpenStudy (amistre64):

pythag it

OpenStudy (amistre64):

144+16 = 160 sqrt(160)

OpenStudy (amistre64):

gotta read the entire q i spose lol

OpenStudy (amistre64):

12^2 + (l+4)^2 = l^2

OpenStudy (amistre64):

got those backwards....

OpenStudy (amistre64):

144 + l^2 = l^2 +16 +8l

OpenStudy (amistre64):

128 -8l = 0 8(16 -l) l = 16; so ladder = 20?

OpenStudy (amistre64):

12^2 +16^2 = 20^2 maybe?

OpenStudy (anonymous):

20 ft looks right to me

OpenStudy (anonymous):

thanks.

OpenStudy (amistre64):

144 256 ---- 400 yep

OpenStudy (anonymous):

medal